OpenStudy (anonymous):

two vectors A and B have components (0,1) and (-1,3), respectively. What is the magnitude of the sum of these two vectors? show work

OpenStudy (amistre64):

what do you get as a new vector after you add them?

OpenStudy (amistre64):

A (0,1) + B (-1,3) --------- (-1,4) ; is what i get

OpenStudy (amistre64):

magnitude is just the length of the vector. which means we pythag it now

OpenStudy (amistre64):

\[\sqrt{1^2+4^2}=\sqrt{17}\]

OpenStudy (amistre64):

do you agree?

OpenStudy (anonymous):

yes it is one of the answers. thank you

OpenStudy (amistre64):

youre welcome